Quincy Quarterback is doing a NFL tryout. He is trying to impress the observers

with his magnificent arm. He can generate a maximum resultant velocity of 19

m/s with his arm. Determine the optimum release angle to maximize throwing

distance. Assume that the ball is release 2.1 m above the ground, and that the ball

hits the ground (it is not caught).


1
Expert's answer
2021-01-26T08:44:34-0500

the movement in the action is the motion of an object launched at a horizontal angle.

we have following quantities.

"v_o = 19\frac{m}{s} \space \space h_m= 2.1m \space \space \theta - ?"

"h_m = \large\frac{v_o^2sin^2\theta }{2g} \to" "sin\theta = \sqrt{\large\frac{2gh_m}{v_o^2}} = \sqrt{\large\frac{2*10*2.1}{361}}=0.3244"

"\theta = 18.93^o"